### v1

# Isentropic processes for axisymmetric Black Holes

Abstract
We demonstrate that the isentropic absorption of a classical charged test particle is classically forbidden for all 3+1 dimensional stationary, non-extremal, axisymmetric black holes in any diffeomorphism invariant theory of gravity. This result is derived purely from the near-horizon geometry and thermodynamic properties of the black hole spacetime, independent of the specific gravitational theory. We further consider the Kerr–Newman black hole in general relativity and analyze within the semiclassical WKB approximation, the conditions under which isentropic absorption may become allowed. Broader implications for the second law and extremality bounds are discussed.
1 Introduction
Beginning in the 1970s, several independent universal entropy bounds were proposed for arbitrary physical systems. Among the earliest was the Bekenstein bound, formulated using semiclassical arguments to suggest that the maximum entropy of an object with finite energy and spatial extent is bounded from above [1, 2, 3, 4]. One can also interpret it as an upper bound on the information contained by a system with a finite energy and spatial extent [2]. In this context, [5] examined the constraints on a channel’s communication capacity imposed by the Bekenstein bound. Further, [6] explored the validity of the generalized Bekenstein bound for the bound states of interacting quantum fields and found that, if interactions are properly included, the entropy bound holds valid. [7] used the unitarity of scattering amplitudes to show that for a self-sustained quantum field object of size $R$, the universal entropy bound can be expressed as $S\leq 1/\alpha$, where the running coupling $\alpha$ is evaluated at the scale $1/R$. Black holes are known to saturate the entropy bound with their entropy expressible as $S=1/\alpha$ where $\alpha=G/R^{d-2}$, and $G$ is Newton’s gravitational constant in dimensions $d$ [7, 8]. Various other derivations and formulations of the bound have been proposed in the literature. For instance, [9] showed from a holographic perspective that the Bekenstein bound holds for any strength of the gravitational coupling under certain assumptions. There are also several alternate or generalized forms of the entropy bounds, such as the causal entropy bound [10], the Unruh-Wald bound [11, 12], generalized Bekenstein bound [13], covariant entropy bound [14, 15], the generalized covariant entropy bound [16, 17], etc.
There are also several interpretations of the Bekenstein bound. For instance [18] illustrated that the Bekenstein bound naturally arises as a consequence of the Pauli exclusion principle when one considers the fundamental degrees of freedom of the black hole to be finite in number and fermionic in nature. Detailed study of the Bekenstein bound for the black holes’ fermionic degrees of freedom can be found in [19, 20]. The Bekenstein bound imposes significant constraints on the information content and thermodynamic behavior of a black hole over its lifetime. [21] applied the Bekenstein entropy bound to analyze the re-distribution of the fundamental degrees of freedom of generic physical systems and found an average information loss due to the entanglement between geometry and fields. Further, [22] used the entropy bound for constraining the information capacity of black hole remnants and concluded that the black hole remnants, as a final stage of black hole evaporation, might not be a viable resolution of the information loss paradox. The energy density of a quantum field can be both positive and negative, provided that the negative energy density in some region of space is balanced by a positive energy density in other regions. [23] illustrated that the Bekenstein bound requires the region containing negative energy density not to be arbitrarily far from the region with positive energy density; that is, there must be a localization of negative energy. One can refer [24, 25, 26] for the consequences of the negative energy density in the time evolution of a black hole.
The black hole information loss paradox suggests that certain fundamental principles of modern physics, namely, unitarity, the applicability of local quantum field theory in describing black hole evaporation, general covariance, the Bekenstein-Hawking entropy relation with the Boltzman entropy of the black hole, and the existence of a semi-classical observer capable of retrieving information from Hawking radiation cannot all simultaneously hold [27, 28, 29, 30]. Further, to resolve such issues, one perhaps needs to take non-perturbative effects into account. These non-perturbative effects may be less dominant than the semiclassical contributions during the initial stages of evaporation; however they can accumulate and provide a significant contribution over the lifetime of the black hole. The relevance of these effects depends upon several factors such as the dimensions of spacetime [31], the memory burden effect [32, 33], fluctuations in the gravitational degrees of freedom [34], existence of critical regimes and instabilities [33, 35], etc. The identification that the Bekenstein-Hawking entropy of a black hole is proportional to its Boltzmann entropy [2] relies on the semiclassical approximation. However, it has been argued in [28] that the non-perturbative effects can lead to the breakdown of the semiclassical approximation and can allow the Boltzmann entropy inside a black hole to be greater than its Bekenstein-Hawking entropy. In particular, [28] analyzed the isentropic absorption of charged particles by a Reisner-Nordstrom black hole within WKB approximation and proposed that such a process could be a non-perturbative way through which entropy bound can be exceeded. In this article, we consider the isentropic absorption of charged particles by a stationary axi-symmetric black hole within WKB approximation processes and discuss implications in the context of the entropy bound.
In section 2, we discuss a general stationary nonextremal axisymmetric black hole spacetime and show that classically it is not possible to absorb a charged particle without changing the black hole’s entropy. There exists a classically forbidden region outside the black hole for such an absorption. However, quantum mechanically, it is possible for certain values of the test particle and the black hole parameters that the test particle tunnel through this potential barrier and get absorbed by the black hole. In section 3 we use the WKB approximation to compute the tunneling probability. Finally in section 4 we discuss the implications of this type of absorption. We use units where $\hbar=c=k_{B}=G=1$, and adopt a mostly positive metric signature ($+$)(-$+$)( - + + + ).
2 Charged particle in stationary axisymmetric spacetime
We consider the following general stationary axisymmetric black hole in 3+1 dimensions [36, 37]:
$ds^{2}=g_{tt}(r,\theta)dt^{2}+2g_{t\phi}(r,\theta)dtd\phi+g_{\phi\phi}(r,\theta)d\phi^{2}+g_{rr}(r,\theta)dr^{2}+g_{\theta\theta}(r,\theta)d\theta^{2}.$ (1)
The metric functions are assumed to be smooth, continuous, atleast $C^{2}$ differentiable, independent of the $t$ and $\phi$ coordinates and thus adapted to the Killing symmetries of stationarity, $\partial_{t}$ and axi-symmetry $\partial_{\phi}$. The spacetime is assumed to be circular and asymptotically flat. Further, the black hole spacetime is assumed to possess a Killing horizon at $r=r_{+}$ such that $g^{rr}(r_{+},\theta)=0$. The black hole is described by its mass $M$, charge $Q$ and angular momemtum $J$.
The trajectory of a particle of mass $m$ and charge $q$ can be described by the following Lagrangian :
$\mathcal{L}=\frac{1}{2}mg_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}+qA_{\mu}\dot{x}^{\mu},$ (2)
where $A_{\mu}=(-A_{t}(r,\theta),0,0,A_{\phi}(r,\theta))$ is the electromagnetic 4-vector potential arising from the stationary and axisymmetric blackhole [38]. Since $t$ and $\phi$ coordinates are cyclic, the corresponding canonical momenta, say $E$ and $L$ respectively, will be conserved. They are given by  [^1] At some places in literature, the canonical momentum conjugate to the cyclic coordinate $\phi$ is denoted by $L_{z}$. We write $L_{z}=L$.
$E\equiv-\frac{\partial\mathcal{L}}{\partial\dot{t}}=-mg_{tt}\dot{t}-mg_{t\phi}\dot{\phi}+qA_{t}$ (3)
$L\equiv\frac{\partial\mathcal{L}}{\partial\dot{\phi}}=mg_{\phi t}\dot{t}+mg_{\phi\phi}\dot{\phi}+qA_{\phi}$ (4)
where dot denotes the derivative with respect to an affine parameter, say $\lambda$. Solving the above equation (3)-(4), one gets the following equation of motion:
$\dot{t}=\frac{E^{\prime}g_{\phi\phi}+L^{\prime}g_{t\phi}}{mD}$ (5)
$\dot{\phi}=-\frac{E^{\prime}g_{\phi t}+L^{\prime}g_{tt}}{mD}$ (6)
where, $D\equiv g_{t\phi}^{2}-g_{tt}g_{\phi\phi}$, $E^{\prime}\equiv E-qA_{t}$, $L^{\prime}\equiv L-qA_{\phi}$. Now, for the classical test particle motion, one can rewrite Eq.(1) as
$-\epsilon=g_{tt}(r,\theta)\bigg{(}\frac{dt}{d\lambda}\bigg{)}^{2}+2g_{t\phi}(r,\theta)\frac{dt}{d\lambda}\frac{d\phi}{d\lambda}+g_{\phi\phi}(r,\theta)\bigg{(}\frac{d\phi}{d\lambda}\bigg{)}^{2}+g_{rr}(r,\theta)\bigg{(}\frac{dr}{d\lambda}\bigg{)}^{2}+g_{\theta\theta}(r,\theta)\bigg{(}\frac{d\theta}{d\lambda}\bigg{)}^{2},$ (7)
where $\epsilon$ is 1 for a massive particle. Taking $\epsilon$ to the right-hand side and using equations (5)-(6), one gets
$\bigg{(}\frac{dr}{d\lambda}\bigg{)}^{2}+g^{rr}(r,\theta)g_{\theta\theta}(r,\theta)\bigg{(}\frac{d\theta}{d\lambda}\bigg{)}^{2}+V_{eff}(r,\theta)=0,$ (8)
where
$V_{eff}(r,\theta)=-\frac{g^{rr}(E^{\prime 2}g_{\phi\phi}+L^{\prime 2}g_{tt}+2g_{t\phi}L^{\prime}E^{\prime})}{m^{2}D}+\epsilon g^{rr}.$ (9)
The first and second terms in equation (8) are positive which implies that the third term denoted by $V_{eff}$ must be negative for the expression to be valid for a classical test particle. Further, we stress that Eqs.(8)-(9) are valid for any diffeomorphism invariant theory of gravity that admits a black hole solution of the form described in (1).
2.1 Isentropic absorption of the charged particle
For an isentropic process, the second law of black hole mechanics leads to
$TdS=0=dM-\Phi_{e}dQ-\Omega dJ.$ (10)
The change in mass $dM$, charge $dQ$, and angular momentum $dJ$ of the blackhole, when the charged particle enters the horizon, is given by $E$, $q$, and $L$ respectively. Further, the change of electromagnetic potential energy is given by $q\Phi_{e}=qA_{t}-\Omega qA_{\phi}$. Therefore, the above equation gives
$E-qA_{t}+q\Omega A_{\phi}-\Omega L=0,$ (11)
which is $E^{\prime}=\Omega L^{\prime}$, at the event horizon. Therefore, the effective potential at the horizon can be written as
$V_{eff}(r=r_{+},\theta)=-\frac{g^{rr}L^{\prime 2}(\Omega^{2}g_{\phi\phi}+g_{tt}+2g_{t\phi}\Omega)}{m^{2}D}\bigg{|}_{r=r+}+\epsilon g^{rr}\bigg{|}_{r=r+}$ (12)
The spacetime under consideration is stationary, so the event horizon will be a Killing horizon [39, 40]. One can consider the Killing horizon to be generated by the Killing vector
$\chi^{\mu}=\partial_{t}+\Omega_{H}\partial_{\phi}.$ (13)
The effective potential at the horizon in terms of $\chi$ can be rewritten as
$V_{eff}(r=r_{+},\theta)=-\frac{g^{rr}L^{\prime 2}\chi^{\mu}\chi_{\mu}}{m^{2}D}\bigg{|}_{r=r+}+\epsilon g^{rr}\bigg{|}_{r=r+}.$ (14)
Figure 1: The above plots show the effective potential experienced by a charged particle for isentropic absorption by a Kerr-Newman blackhole in general relativity. The expression for the effective potential is given in Eq.(9), with $E^{\prime}(r)\equiv E-qA_{t}(r)=qA_{t}(r_{+})+\Omega(L-qA_{\phi}(r_{+}))-qA_{t}(r)$ for the isentropic absorption. The unit of $\theta$ is taken to be radians, so the plots are near the equator.
The effective potential given above in Eq.(14) is independent of the sign of $L^{\prime}$. One can use the property of axisymmetry of the spacetime and the zeroth law of black hole mechanics to show that $\chi^{\mu}\chi_{\mu}$, $g^{rr}$, and $D$ all vanish linearly at the Killing horizon of a stationary axisymmetric non-extremal black hole [41], i.e., as follows:
$\chi^{\mu}\chi_{\mu}\approx-2\kappa(r-r_{+})+\mathcal{O}((r-r_{+})^{2});D\approx 2\kappa(r-r_{+})g_{\phi\phi}+\mathcal{O}((r-r_{+})^{2});g^{rr}\propto(r-r_{+}).$ (15)
Here, $\kappa$ is the surface gravity which is constant over the Killing horizon from the zeroth law of black hole mechanics  [^2] One can refer to [42] for a proof of the zeroth law of black hole mechanics that does not depend on any assumptions about the dynamics of the theory, but only assumes the existence of a Killing horizon with a nonzero $\kappa$.. Hence, the above effective potential $V_{eff}(r_{+},\theta)$ vanishes at the horizon for an isentropic absorption of the test charged particle. Taking the derivative of Eq.(9) with respect to $r$ one gets after some manipulation:
$\frac{dV_{eff}}{dr}=g^{\prime rr}g_{rr}V_{eff}-\frac{D^{\prime}}{D}(V_{eff}-g^{rr}\epsilon)-\frac{g^{rr}}{m^{2}D}(E^{\prime 2}g^{\prime}_{\phi\phi}+L^{\prime 2}g^{\prime}_{tt}+2L^{\prime}E^{\prime}g^{\prime}_{t\phi}),$ (16)
where prime on $D$ and the metric components denote the derivative with respect to $r$. Evaluating the above expression on the horizon for an isentropic process with $E^{\prime}=\Omega L^{\prime}$ one gets,
$\frac{dV_{eff}}{dr}\bigg{|}_{r=r_{+}}=\frac{\epsilon D^{\prime}g^{rr}}{D}\bigg{|}_{r=r_{+}}+\frac{L^{\prime 2}g^{rr}}{m^{2}D}\bigg{[}\frac{D^{\prime}\chi^{\mu}\chi_{\mu}}{D}-\frac{g^{\prime rr}\chi^{\mu}\chi_{\mu}}{g^{rr}}-(\chi^{\mu}\chi_{\mu})_{,r}\bigg{]}\bigg{|}_{r=r_{+}}$ (17)
Now, using L’Hospital rule for taking limit near horizon and the property that $D$, $\chi^{\mu}\chi_{\mu}$, and $g^{rr}$ all go to zero near horizon linearly as shown in Eq.(15), for the nonextremal case, one gets
$\frac{dV_{eff}}{dr}\bigg{|}_{r=r_{+}}=g^{\prime rr}\bigg{[}\epsilon-\frac{L^{\prime 2}(\chi^{\mu}\chi_{\mu})_{,r}}{m^{2}D^{\prime}}\bigg{]}_{r=r_{+}}=g^{\prime rr}\bigg{[}\epsilon+\frac{L^{\prime 2}}{m^{2}g_{\phi\phi}}\bigg{]}_{r=r_{+}}>0,$ (18)
since $g^{\prime rr}>0$ and $g_{\phi\phi}>0$ at the horizon. Since $V_{eff}(r_{+},\theta)=0$ and $dV_{eff}(r_{+},\theta)/dr>0$, we necessarily have $V_{eff}>0$ outside the horizon and hence, classically it is not possible to absorb the charged particle isentropically.
At a large distance from the black hole, the electromagnetic 4-vector potential vanishes, i.e, $A_{t}(r\rightarrow\infty,\theta)\rightarrow 0,A_{\phi}(r\rightarrow\infty,\theta)\rightarrow 0$. Therefore, $E^{\prime}\rightarrow E,L^{\prime}\rightarrow L$ at spatial infinity. Further, $g^{rr}\rightarrow 1$, $g_{tt}\rightarrow-1,g_{t\phi}\rightarrow 0,g_{\phi\phi}\rightarrow\infty$ at spatial infinity. So, the effective potential shown in Eq.(9) takes the following form:
$V_{eff}(r\rightarrow\infty,\theta)\rightarrow-\frac{E^{2}}{m^{2}}+\epsilon.$ (19)
The above asymptotic expression for the effective potential is consistent with that expected for a flat spacetime. For a particle at rest at infinity, Eq. (19) gives $V_{\text{eff}}(r\rightarrow\infty,\theta)\rightarrow 0$, while it becomes negative for a moving particle. For the isentropic absorption, the conserved energy $E=\Omega L+qA_{t}(r=r+,\theta)-\Omega qA_{\phi}(r=r+,\theta)>m^{2}$, which is possible for an appropriate choice of parameters $\{q,m,L,Q,M,a\}$. So, the effective potential at infinity is negative for this set $\{q,m,L,Q,M,a\}$, which implies that there also exists a classically allowed region for $r>r_{2}$ for some $r_{2}>r_{+}$ apart from the classically forbidden region for $r_{+}<r<r_{2}$.
2.2 Kerr-Newman black holes in general relativity
So far, our discussion regarding the isentropic absorption of a charged particle by a stationary axisymmetric nonextremal black hole was quite general. For the sake of definiteness, we now restrict to the non-extremal Kerr-Newman black hole in general relativity  [^3] Since the surface gravity, $\kappa$, is zero in the extremal case, the discussion in the previous section may not be true for the extremal Kerr-Newmann black hole.. In general relativity, the Kerr-Newman black hole in 3+1 dimensions, in Boyer–Lindquist coordinates, is described by the metric [36, 38]
$ds^{2}=$ $-\frac{\Delta-a^{2}\sin^{2}\theta}{\Sigma}\,dt^{2}-\frac{2a\sin^{2}\theta(r^{2}+a^{2}-\Delta)}{\Sigma}\,dt\,d\phi+\frac{(r^{2}+a^{2})^{2}-\Delta a^{2}\sin^{2}\theta}{\Sigma}\sin^{2}\theta\,d\phi^{2}$
$+\frac{\Sigma}{\Delta}\,dr^{2}+\Sigma\,d\theta^{2},$ (20)
with
$\Sigma(r,\theta)$ $=r^{2}+a^{2}\cos^{2}\theta,$ (21)
$\Delta(r)$ $=r^{2}-2Mr+a^{2}+Q^{2}.$ (22)
Here, $a$ is the specific angular momentum, $M>0$ is the ADM mass, and $Q$ is the electric charge of the blackhole. Assuming a vanishing magnetic charge, the electromagnetic 4-vector potential of the black hole is given by $A_{\mu}=(-Qr/\Sigma,0,0,Qar\sin^{2}\theta/\Sigma)$. So, the sign convention is such that a positive value of $Q$ means a positively charged black hole. One can see by substituting the metric Eq.(2.2) in equations (12) and (16) that for the isentropic absorption of the charged particle
$V_{eff}\bigg{|}_{r=r_{+}}$ $=0$ (23)
$\frac{dV_{eff}}{dr}\bigg{|}_{r=r_{+}}$ $=\frac{2(Mr_{+}-a^{2}-Q^{2})}{r_{+}^{3}}\bigg{[}1+\frac{L^{\prime 2}(r_{+}^{2}+a^{2}\cos^{2}\theta)}{(r_{+}^{2}+a^{2})^{2}}\bigg{]}>0.$ (24)
by direct calculation. Therefore, the isentropic absorption of the charged test particle is not possible classically.
Figure 2: The above plots show the Euclidean action experienced by a charged particle for isentropic absorption by a Kerr-Newman blackhole in general relativity. The turning points exist only for certain values of parameters. The plots used the parameter values for which the turning point exists.
Figure 3: The above plots show the Euclidean action experienced by a charged particle for isentropic absorption vs the radius of the outer horizon of a Kerr-Newman blackhole in general relativity. The turning points exist only for certain values of parameters. The plots used the parameter values for which the turning points exist.
3 Quantum tunneling
We saw in the previous section that the effective potential vanishes at the event horizon and that its derivative becomes positive for the isentropic absorption. We plot the effective potential in Fig.[1], from which it can be seen that for certain values of charge, mass, and angular momentum, the effective potential can become negative again at some distance from the event horizon. The region between the event horizon at $r=r_{+}$ and the point outside, say $r=r_{2}>r_{+}$, where the effective potential vanishes again, is classically forbidden. We see from Fig. [1] that increasing the angular momentum of the particle increases both the width of the classically forbidden region and the height of $V_{eff}$. However, the increase in charge reduces both. Further, we see that the change of mass of the test particle doesn’t change the width of the classically forbidden region, but the height of the potential barrier experienced changes. So, the massive charged test particles with low angular momentum experience a small, but nonzero, potential for the isentropic absorption.
The tunneling probability through the classically forbidden region between $r_{+}$ and $r_{2}$ in the WKB approximation is given by [28]
$\Gamma\approx e^{-2S_{E}},$ (25)
where $S_{E}$ is the Euclidean action given by
[…]
Here, we have assumed $d\theta/d\lambda=0$ for simplicity and used Eq.(8) with $\lambda_{E}=i\lambda$. We choose negative root in $dr/d\lambda_{E}=\pm\sqrt{V_{eff}}$ because the particle is being absorved by the black hole. It’s difficult to evaluate the integral in Eq. (29) analytically, so we compute it numerically and plot the corresponding Euclidean action for a real value of turning points $r_{+}$ and $r_{2}$. The tunneling probability in Eq.(25) decreases (increases) as the Euclidean action increases (decreases), so Fig.[2]-[3] effectively describe the tunneling probability in the WKB approximation.
One can notice from Fig.[2] that the Euclidean action increases with an increase in the angular momentum of the particle for all combinations of parameters. Therefore, the tunneling probability for low angular momentum particles is large in comparison with the higher angular momentum particles. One can understand it as the effective potential as well as the width of the classically forbidden region also increases with an increase in the angular momentum of the test particle. Furthermore, the increase in the magnitude of charge $Q$ of the blackhole also reduces the Euclidean action. However, the increase in mass of the blackhole increases the Euclidean action. The increase of the specific angular momentum $a$ of the black hole decreases the effective action for the positive $L$ particles, but increases the effective action for the negative $L$ particles. Therefore, a low mass, high specific angular momentum, and large charge black hole has a higher tunneling probability for a test charged particle. The Euclidean action is smaller for a larger polar angle. Therefore, the isentropic absorption is more probable near the equator of the black hole.
In Fig.[3], we see that the increase of the size of the black hole, determined by its outer horizon $r_{+}$, increases the Euclidean action for all combinations of parameters. Therefore, smaller black holes possess a greater tunneling probability than larger black holes. This suggests that the smaller black holes are relatively less classical as they can absorb charged particles isentropically and are prone to violate the entropy bound more efficiently than a large black hole. Furthermore, we notice that increasing the mass and the magnitude of angular momentum of the test particle increases the Euclidean action for a fixed size $r_{+}$ of the black hole, whereas an increase in the magnitude of the charge of the test particles decreases the Euclidean action. Therefore, lighter particles with small angular momentum, but large charge, have a large tunneling probability. Further, the tunneling probability is larger near the equator for a fixed $r_{+}$.
4 Discussion
We investigated whether a stationary axisymmetric black hole can absorb a classical test charged particle without a change in the black hole’s entropy, and found that this process is classically forbidden in any diffeomorphism-invariant theory of gravity. Further, we analysed a Kerr-Newman blackhole as a specific example. A test charged particle outside a Kerr–Newman black hole experiences an effective potential for isentropic absorption, which remains positive between the outer horizon and a certain distance away from it. So, classically, this region outside the black hole is forbidden. However, for a certain range of mass, charge, and angular momentum of a test charged particle, it is quantum mechanically possible to tunnel through this classically forbidden region.
Our results show that the tunneling probability increases with a decrease in mass of either the test particle or the black hole. However, the tunneling probability increases with an increase in the charge of either the test particle or the black hole. Furthermore, we also find that the tunneling probability is large for a small blackhole in comparison to the larger black holes. These observations suggest that the low mass and high charge black holes with a smaller radius of the outer horizon are more prone to isentropic absorption of classical test charged particles. Moreover, if a charged particle that is entangled with degrees of freedom outside the black hole is isentropically thrown into the black hole, the entanglement between the inside and outside of the black hole increases without any change in the areal entropy. As emphasized in [28], this raises important questions about whether the black hole can accumulate entanglement entropy isentropically over time. Since the Boltzmann entropy accounts for the total number of accessible microstates and is typically greater than the entanglement entropy [28, 43, 44, 45], could it perhaps happen through such a process that the Boltzmann entropy becomes greater than the Bekenstein-Hawking entropy?